Todd is 12 years older than Delores. In 2 years, Todd will be twice as old as Delores. How old is Todd now?
Answer by awang1996(50)

d=Delores
t=Todd
d+12=t
2(d+2)=t+2
2d+4=t+2
2d+4=d+12+2
2d+4=d+14
Subtract d from both sides.
d+4=14
Subtract 4 from both sides.
d=10
t=22